OpenStudy (anonymous):

Implicit function: cos(6xu) = 2u + 7x

OpenStudy (anonymous):

a) Calculate the derivative of u with respect to x. u ' (x) = ?? b) Calculate the derivative of x with respect to u. x ' (u) = ??

hartnn (hartnn):

u tried? what u got ?

OpenStudy (anonymous):

for u' i had: \[u'=\frac{ -7\sin(6xu) }{ 2 }\]

OpenStudy (anonymous):

which def is not right.

hartnn (hartnn):

cos(6xu) = 2u + 7x -sin(6xu) (6xu)' = 2u' +7

hartnn (hartnn):

apply product rule for (6xu)'

OpenStudy (anonymous):

would it by 6x * u = (6x)u'+6u

hartnn (hartnn):

yup. now isolate u'

OpenStudy (anonymous):

do i take my 2u' over to the right hand side?

hartnn (hartnn):

yes

OpenStudy (anonymous):

i had: \[u'=\frac{ 2u'+7 }{ -\sin(6xu)(6x) }\]

OpenStudy (anonymous):

so then it's u'-2u'= what i just wrote (minus the 2u' on the top)

OpenStudy (anonymous):

so it's \[-u'=\frac{ -7 }{ \sin(6xu)(6x) }\]

hartnn (hartnn):

u didn't isolate u'

OpenStudy (anonymous):

i thought i did?

hartnn (hartnn):

-sin(6xu) ( (6x)u'+6u) = 2u' +7 -sin(6xu) (6x)u' + 6u (-sin(6xu)) -2u' =7 {-sin(6xu) (6x)-2}u'=7- 6u (-sin(6xu)) u' = {7+ 6u (sin(6xu))} / {-sin(6xu) (6x)-2} just some algebraic manipulations

OpenStudy (anonymous):

these are my possible answers, i dont seem to see that one there even though it makes sense to me what you did.

OpenStudy (anonymous):

oh wait, those are the answers for part b. hld one one sec

OpenStudy (anonymous):

those are the possible answers for a

hartnn (hartnn):

those are lots of options....

OpenStudy (anonymous):

the second attachment is the list of possible answers.

OpenStudy (anonymous):

for a

hartnn (hartnn):

i can see 4th option matching with our answer.....

hartnn (hartnn):

{-sin(6xu) (6x)-2}= -2{ 1+3x sin(6xu)}

OpenStudy (anonymous):

they just have stuff factored out and they put the 1/2 on the outside of the equation. i see.

hartnn (hartnn):

did u get all steps ??

OpenStudy (anonymous):

yup! i wrote them all down. : ) so now i just do the same thing for part b but, i do it for x' and not u' ?

hartnn (hartnn):

yeah. u need to diff. again, now w.r.t u

OpenStudy (anonymous):

im still getting u' when i do out my derivative.

hartnn (hartnn):

derivative of u =1

OpenStudy (anonymous):

should i write down x as x' and not one?

hartnn (hartnn):

yup, here x' is dx/du and not dx/dx

OpenStudy (anonymous):

so now i have: \[-\sin(6xu)(6x)(6x')(u)=2+7x'\]

hartnn (hartnn):

cos(6xu) ' = -sin (6xu) (6xu)' = -sin (6xu) (6x+6ux')

OpenStudy (anonymous):

right, i just wrote it down weird lol.

OpenStudy (anonymous):

so that =2+7x'

hartnn (hartnn):

yup, now algebraic manipulations.....

OpenStudy (anonymous):

-sin(6ux)(6ux'+6x)(-sin(6xu)-7x'=2 ???

OpenStudy (anonymous):

prob not .. lol, this is the part i find the hardest because there's so many variables.

hartnn (hartnn):

let a=-sin(6ux) (just temporarily) a(6ux'+6x) = 2+7x now try

hartnn (hartnn):

a(6ux'+6x) = 2+7x'

OpenStudy (anonymous):

a(6ux')+a(6x)=2+7x'

OpenStudy (anonymous):

so... -sin(6ux')-sin(6x)=2+7x'

hartnn (hartnn):

u still didn't isolate x'

hartnn (hartnn):

a(6ux')+a(6x)=2+7x' bring 7x' to left and 6ax to right

hartnn (hartnn):

then factor x' from left

hartnn (hartnn):

need help with that ?

OpenStudy (anonymous):

x'[a6u-7]=a6x+2

hartnn (hartnn):

x'[a6u-7]=-a6x+2

OpenStudy (anonymous):

thats what i said lol.

OpenStudy (anonymous):

so then i sub back in -sin(6ux)

OpenStudy (anonymous):

im not seeing my answer on the list of answers tho ..

hartnn (hartnn):

u just missed a minus.... x'[a6u-7]=-a6x+2 x' = {-a6x+2} / {[a6u-7]} now put a=-.....

OpenStudy (anonymous):

sin(6ux)(6x)+2/-sin(6ux)(6u)-7

hartnn (hartnn):

2nd option

hartnn (hartnn):

factor 2 from num., and - from denom.

OpenStudy (anonymous):

ohh okY I see!!!

hartnn (hartnn):

:)

OpenStudy (anonymous):

thanks a lot, you're really helpful!

hartnn (hartnn):

welcome ^_^